# Timing characterization of MALTA and MALTA2 pixel detectors using Micro X-ray source

### v1

@@ 6.2 Measurements using MALTA2 sample
Figure 9: Delay between the input and output signals using 12kV X-rays on the MALTA2 sample. The contribution to the tail part comes from the corner pixel hits, front-end electronics as well as due to comparison errors from the oscilloscope.
While the set-up for the MALTA2 chip mirrors that of MALTA, distinctions in the MALTA2 configuration necessitates corresponding alterations in connection and firmware. Specifically, the default FPGA firmware for MALTA2 is changed accordingly to receive signals from the computer, subsequently triggering the laser and capturing the timing information when X-rays interact with the laser. These firmware modifications are crucial to facilitate the transmission of signals to the oscilloscope, ensuring data collection and analysis. The timing study of the MALTA2, Epitaxial chip with an extra deep p-well implant is done over a several range of high voltages with different threshold values. The delay for 12kV voltage can be seen in Figure 9 below.
@@ 7 Results
Figure 10: Comparison between RMS of the timing delay of MALTA and MALTA2 samples using the first 25 ns only. The x-axis shows the input voltage provided to the micro X-ray tube, and the y-axis is the standard deviation of the delay in the first 25 seconds. The timing delay contributions from the set-up are included for both the MALTA and MALTA2 samples.
@@
Figure 11: Comparison between the standard deviation of the timing delay of MALTA and MALTA 2 samples using the Gaussian fit method. The peaks are fitted with a single Gaussian, and the width of the Gaussian fit is compared for MALTA and MALTA2 at different X-ray tube voltages. The x-axis shows the input voltage provided to the micro X-ray tube, and the y-axis is the standard deviation of the Gaussian fit. The timing delay contributions from the experimental set-up are included for both the MALTA and MALTA2 samples.
@@
A distinct and extended tail is evident in both the MALTA and MALTA2 datasets, and it is crucial to emphasize that this phenomenon is not solely attributed to timing differences. Rather, the presence of such pronounced tails can be largely attributed to errors introduced by the oscilloscope. A contribution to the tail part also comes from MALTA frond-end electronics. The underlying mechanism involves the acquisition of two distinct signals from the FPGA in the oscilloscope—one originating from the trigger and the other from the MALTA2 output. Notably, these signals are characterized by the presence of multiple peaks rather than a singular peak. Our approach to determining the timing difference involves the selection of corresponding peaks; however, occasional discrepancies result in the acceptance of the timing associated with the subsequent peak rather than the intended one, leading to the augmentation of data points in the far-tail region. In response to this effect, we adopted two distinct methodologies. Initially, we sought to ascertain the root mean square (RMS) value by excluding the prolonged tail and focusing solely on the initial 25 ns (bunch crossing time for LHC), as shown in Figure 10. From this plot, it is evident that the RMS is increasing with the applied voltages to the X-ray tube, which is solely due to the involvement of a higher number of transition lines in Cu and Cr at higher applied voltage.
@@
Subsequently, an alternative approach involved a comparative analysis of the peak width for both MALTA2 and MALTA datasets. Despite the fact that these peaks do not match perfectly to Gaussian fits, for the sake of general applicability, our analysis considered only single Gaussian fits to derive approximate values. Then, we compared the width values of MALTA and MALTA2 from the fit at different applied voltages to the X-ray tube.
@@ 8 Conclusion
The study was focused on investigating the deviation in the delay time of MALTA and MALTA2 epitaxial samples. Given the limitations of conventional laser techniques, we employed a micro X-ray source comprising copper and chromium layers to generate X-rays. The utilization of a Low-Gain Avalanche Detector (LGAD) allowed for a comprehensive exploration of the properties of the Micro-X-ray source, and we observed that the $k_{\alpha}$ energy levels were consistent with theoretical values. The findings revealed that the micro X-ray source contributed approximately 300 ps to the overall delay time for different laser frequencies, as determined through laser source experiments. Also, the FPGA is found to have a contribution of approximately 80 ps, resulting in the RMS of the delay from MALTA and MALTA2 EPI samples to be approximately 2.8 ns and 2.1 ns with a best value of 2.7 ns and 2 ns, respectively. Also, this experiment concluded the delay time improvement for the second generation of MALTA pixel detectors than its previous generation, providing strong support for its usability in future collider experiments.
@@ Acknowledgments
This project has received funding from the European Unions Horizon 2020 Research and Innovation programme under Grant Agreement numbers: 101004761 (AIDAinnova), 675587 (STREAM), 654168 (AIDA-2020). Also, we thank Science and Engineering Research Board, India and the Ministry of Human Resource Development, India for their funding support.

### v2

@@ 6.2 Measurements Using MALTA2 Sample
Figure 9: Delay between the input and output signals using 12 kV X-rays on the MALTA2 sample. Contributions to the tail part come from corner pixel hits, front-end electronics, as well as comparison errors from the oscilloscope.
While the set-up for the MALTA2 chip mirrors that of MALTA, distinctions in the MALTA2 configuration necessitate corresponding alterations in connection and firmware. Specifically, the default FPGA firmware for MALTA2 is modified to receive signals from the computer, subsequently triggering the laser and capturing the timing information when X-rays interact with the sensor. These firmware modifications are crucial to facilitate the transmission of signals to the oscilloscope, ensuring proper data collection and analysis. The timing study of MALTA2, an epitaxial chip with an extra deep p-well implant, is conducted over a range of high voltages with different threshold values. The delay for 12 kV operation can be seen in Figure 9.
@@ 7 Results
Figure 10: Comparison of the timing delay RMS for MALTA and MALTA2 samples, evaluated within the first 25 ns. Subfigures in (a) show the timing distribution for MALTA, and Subfigures in (b) show the timing distribution for MALTA2. The x-axis represents the input voltage applied to the micro X-ray tube, while the y-axis indicates the number of hits within the first 25 ns. The mean and RMS values of the respective distributions are displayed in each panel. Timing delay contributions from the measurement set-up are included for both MALTA and MALTA2 samples.
@@
Figure 11: Comparison between the standard deviation of the timing delay of MALTA and MALTA2 samples using the Gaussian fit method. The peaks are fitted with a single Gaussian, and the width of the Gaussian fit is compared for MALTA and MALTA2 at different X-ray tube voltages. The x-axis shows the input voltage applied to the micro X-ray tube, and the y-axis represents the standard deviation of the Gaussian fit. Timing delay contributions from the experimental set-up are included for both the MALTA and MALTA2 samples.
@@
A distinct and extended tail is evident in both the MALTA and MALTA2 datasets. It is crucial to emphasize that this phenomenon is not solely attributed to timing differences. Rather, the presence of such pronounced tails can largely be attributed to errors introduced by the oscilloscope. Additional contributions to the tail arise from the MALTA front-end electronics. The underlying mechanism involves the acquisition of two distinct signals from the FPGA in the oscilloscope—one originating from the trigger and the other from the MALTA2 output. These signals are characterized by multiple peaks rather than a single peak. Our approach to determining the timing difference involves selecting the corresponding peaks; however, occasional discrepancies result in the acceptance of the timing associated with the subsequent peak rather than the intended one, leading to the augmentation of data points in the far-tail region.
To mitigate this effect, we adopted two distinct methodologies. Initially, we calculated the root mean square (RMS) value by excluding the prolonged tail and focusing solely on the initial 25 ns (bunch-crossing time for the LHC), as shown in Figure 10. From this plot, it is evident that the RMS increases with the applied voltage to the X-ray tube, primarily due to the involvement of a higher number of transition lines in Cu and Cr at higher applied voltage.
@@
Subsequently, an alternative approach involved a comparative analysis of the peak width for both MALTA2 and MALTA datasets. Although these peaks do not perfectly match Gaussian fits, our analysis considered only single Gaussian fits for general applicability to derive approximate values. We then compared the width values of MALTA and MALTA2 from the fit at different applied voltages to the X-ray tube.
From Figure 11, we find that the RMS of the peak is approximately 2.6 ns for MALTA2, with a best value of 2.5 ns. This value includes contributions from the setup ($\sim 300$ ps), jitter from the FPGA ($\sim 160$ ps), and jitter from the laser source (up to 50 ps), all of which affect the final measurement. After correcting for these contributions, the RMS of the fast timing is calculated as
$\sqrt{2600^{2}-300^{2}-160^{2}-50^{2}}\ \text{ps}\approx 2570~\text{ps}\ (2.57~\text{ns}),$
with the best measured value of 2.5 ns for the MALTA2 epitaxial sample.
@@
From test-beam data collected between 2021 and 2023, the best timing resolution achieved for MALTA2 with a Czochralski substrate at low threshold settings was 1.7 ns using a cluster size of two pixels [16]. In contrast, our measurements were performed on the full chip without clustering, yet we obtained a comparable timing resolution in laboratory conditions without the need for high-intensity beams. This highlights the practicality of our approach and its potential usefulness for future measurements.
@@ 8 Conclusion
This study demonstrates the feasibility of using a laboratory-based micro X-ray tube for timing measurements, providing a convenient alternative to test beam campaigns. The study focused on investigating the delay time in MALTA and MALTA2 samples. Given the limitations of conventional laser techniques, a micro X-ray source comprising copper and chromium layers was employed to generate X-rays. The use of a Low-Gain Avalanche Detector (LGAD) enabled detailed characterization of the micro X-ray source, where the observed $K_{\alpha}$ energy peaks were consistent with theoretical values. Our findings show that the micro X-ray source contributes approximately 300 ps to the overall delay time, while the FPGA adds $\sim 160$ ps and the laser jitter contributes up to 50 ps. After accounting for these effects, the RMS delay time is found to be $\sim 3.3$ ns (best value of 3.2 ns) for MALTA and $\sim 2.6$ ns (best value of 2.5 ns) for MALTA2. This confirms the improved performance of the second generation of MALTA pixel detectors over the first, providing strong evidence of their suitability for future collider experiments.
@@ Acknowledgments
This project has received funding from the European Unions Horizon 2020 Research and Innovation programme under Grant Agreement numbers: 101004761 (AIDAinnova), 675587 (STREAM), 654168 (AIDA-2020). Also, we thank Science and Engineering Research Board, India, and the Ministry of Human Resource Development, India, for their funding support.